OpenStudy (anonymous):

John and Martha each had a steak dinner last night at Fred's diner. The steaks were $12.95 apiece, and they each had a glass of wine at $6 apiece. The bill included an additional 8% tax. How much was the total of John and Martha's bill?

OpenStudy (jdoe0001):

\(\large \begin{array}{lrlll} steak&12.95\\ &12.95\\ wine&6.00\\ &+6.00\\ \hline\\ subtotal&{\color{red}{ \square }}\\ tax&8\%\to \frac{8}{100}\cdot {\color{red}{ \square }}\to +0.08{\color{red}{ \square }}\\ \hline\\ total&? \end{array}\)

OpenStudy (anonymous):

subtotal $37.90 tax 8% $3.03 Total 40.93